OpenStudy (anonymous):

Ship A and Ship B leave the same port at noon. Ship A is sailing east at 35 km/h and ship B is sailing north at 25 km/h. How fast is the distance between the ships changing at 4:00 p.m.?

OpenStudy (anonymous):

I know this is a related rates problem. and I did: a^2+b^2 = c^2 a= 140, b = 100 c = 172.0465 differentiate the function by each variable and solved for dc/dt and I got 43.0116 = dc/dt Am I correct ?

OpenStudy (anonymous):

yes, no anyone ? This is my first time attempting to do this by myself completely and I just need to know if I'm correct?

OpenStudy (anonymous):

I think your on the right track..it looks right to me but I'm not math expert